OpenStudy (anonymous):

Solve the equation using the Zero-Product Property. (x-2)(x+7)=0

jimthompson5910 (jim_thompson5910):

If (x-2)(x+7)=0, then x-2 = 0 or x+7 = 0

jimthompson5910 (jim_thompson5910):

From here, solve each equation for x

OpenStudy (anonymous):

This rule tells us that a product is equal to zero when at least one of its factors is equal to zero. So x-2=0 => x=2 or x+7=0 => x=-7.